# A Coordinated Routing Approach for Enhancing Bus Timeliness and Travel Efficiency in Mixed-Traffic Environment

### v1

@@ I Introduction
Connected and automated vehicles (CAVs) have emerged as a central element in the evolution of modern transportation systems. Advances in vehicle-to-vehicle, vehicle-to-infrastructure, and vehicle-to-everything communication technologies highlight the significant potential of CAVs in improving traffic efficiency, reducing energy consumption, and enhancing road safety [1, 2]. The adoption of CAVs is projected to substantially increase in the coming decades; the transition towards a fully autonomous fleet will occur gradually [3]. Transportation systems will experience prolonged periods of mixed traffic conditions where CAVs and human-driven vehicles (HVs) coexist. Understanding how to optimize the efficiency under increasing penetration rates of CAVs during this transitional period is critical [4, 5, 6]. Extensive experimental studies have indicated that heterogeneous traffic environments involving CAVs can effectively reduce total travel times [7, 8, 9]. Implementing suitable right-of-way (ROW) management strategies is essential for ensuring the seamless integration of CAVs into existing transportation networks and for fully leveraging their benefits throughout this transitional phase.
@@
To address the tradeoff inherent in implementing DLs for CAVs, a promising strategy has been proposed: the CAV and bus joint DL. Unlike conventional approaches that require constructing new lanes or displacing HVs, joint use of DLs by buses and CAVs repurposes existing DLs, which were originally designed for buses to ensure their on-time arrivals and maintain public transit efficiency [16]. While these lanes provide dedicated driving environments for transit operations, the reserved space cannot taken by other non-transit vehicle lower the traffic efficiency  [17]. Allowing CAVs to access these lanes can enhance lane utilization while maintaining transit reliability. In the existing literature, Luo et al. developed a framework for the optimal placement of CAV DLs to maximize benefits for all vehicles in the system [18]. Chen et al. explored a space-time modeling framework that demonstrates the mixed-use of DLs can enhance CAV travel efficiency and alleviate the overall traffic congestion [19]. A systematic framework investigating a variety of lane policies, such as exclusive automated vehicle lanes, regular vehicle lanes, and mixed-use lanes, were investigated in [11]. While the developed approaches provide new insights into the potential benefits of jointly using DLs, these methods remain limited in their ability to dynamically manage CAVs’ routing, particularly in coordinating vehicle behaviors across the network to balance traffic congestion reduction with maintaining reliable bus travel times, especially under fluctuating demand and a dynamic travel environment.
In this paper, we propose to develop a dynamic routing approach to coordinate CAVs traveling in a mixed traffic network, allowing for the DL sharing with buses. The goal is to prioritize the buses’ adherence to their scheduled timeliness while taking advantage of the joint DLs for CAVs to improve road capacity, reduce congestion, and enhance system-wide efficiency. By leveraging real-time traffic data, the proposed model enables estimation of potential tense flow on network edges and dynamically navigates the rerouting of CAVs to mitigate congestion on DLs. The contributions of this work are summarized as follows: (i) We model the rerouting problem of CAVs in a mixed traffic network where buses and CAVs jointly use DLs to fully take advantage of DLs while enhancing buses’ adherence to schedules. (ii) We incorporate a dynamic route condition estimation scheme in the rerouting framework, enabling real-time monitoring of traffic flow and timely adjustment of CAV routes when potential DL congestion is detected. Simulation performed in SUMO demonstrates the effectiveness of the proposed approach, indicating significant improvements in traffic efficiency through coordinated use of joint DLs.
@@ II Problem Statement
Figure 1: The road transportation network, where bus stations ($\mathcal{S}$) are shown by red nodes, intersections ($\mathcal{I}$) are shown in black, and the dedicated lanes are shown by red edges.
Consider an urban road network represented by a directed graph $G\!=\!(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ denotes the set of nodes and $\mathcal{E}\!\subseteq\!\mathcal{V}\!\times\!\mathcal{V}$ denotes the set of directed edges. Taking the road network of Manhattan as an example, as shown in Fig. 1, the node set $\mathcal{V}$ is comprised of two subsets, denoted as $\mathcal{V}\!=\!\mathcal{S}\cup\mathcal{I}$, where $\mathcal{S}$ represents the set of bus stations and $\mathcal{I}$ represents the set of road intersections. Each directed edge $(v,v^{\prime})\in\mathcal{E}$ models a road segment from node $v$ to node $v^{\prime}$, with $v,v^{\prime}\in\mathcal{V}$. Consider each edge consisting of two lanes in the same direction, i.e., joint DL and GPL, where the GPL can be used by both CAVs and HVs, while DLs are reserved only for joint use by buses and CAVs.
Throughout this paper, we consider three types of vehicles traveling in the transportation network, including HVs, CAVs, and buses. The sets of HVs, CAVs, and buses within the monitoring time horizon are denoted by $\mathcal{N}^{\text{hv}}$, $\mathcal{N}^{\text{cav}}$, and $\mathcal{N}^{b}$, respectively. Accordingly, the set of all vehicles is represented as $\mathcal{N}\!=\!\mathcal{N}^{\text{hv}}\!\cup\!{\mathcal{N}^{\text{cav}}}\!\cup\!{\mathcal{N}^{b}}$. As illustrated in Fig. 2, buses use only DL and aim to arrive at each bus stop on time, following a specific timetable. In contrast, HVs can use only GPLs for travel purposes and route planning. With higher flexibility, CAVs can switch between joint DLs and GPLs to improve travel efficiency while minimizing interruptions to buses. Compared to employing only GPLs, driving on joint DLs enables CAVs to keep a smaller inter-vehicle distance, thus improving road capacity while avoiding disturbance caused by other HVs on GPLs.
@@ III-A Traffic Flow on Joint Dedicated Lane
where $v_{i,\ell}\!\in\!\mathcal{I}$, $\ell\!=\!1,\dots,n_{i}$, denotes the $\ell$th road intersection along the planned route of CAV $i$ toward its destination. To alleviate traffic congestion on DLs that could lead to delays in bus arrivals at their stops, the cumulative flow of CAVs entering each edge on the DLs is actively monitored. Specifically, let $\tau_{i,\ell}^{0}$ represent the free-flow travel time on edge $(v_{i,\ell},v_{i,\ell+1})$, and $t_{i,\ell}$ denote the arrival time of CAV $i$ at intersection $v_{i,\ell}$. Thus, the arrival time of CAV $i$ at the next intersection $v_{i,\ell+1}$ is of the form
@@ III-B Traffic Flow on General-Purpose Lane
where the indicated function $I_{(v,v^{\prime})}(t_{\text{sys}},\Delta\tilde{T})$ characterizes the number of CAVs anticipated to traverse the edge $(v,v^{\prime})$ within the monitoring time window. Mathematically, it is denoted as
@@
The traffic flow monitoring windows are defined by $\Delta{T}$ in (5) for joint DLs and by $\Delta{\tilde{T}}$ in (9) for GPLs, respectively, to capture real-time traffic conditions. A larger monitoring window puts greater emphasis on the macroscopic characteristics of traffic flow on network edges. With different emphasis, the sizes of the two monitoring windows are not necessarily to be identical.
@@ IV-A Identification of CAVs for Rerouting
Let $\hat{t}_{b,k}$ denote the estimated arrival time of bus $b\!\in\!\mathcal{N}^{\text{b}}$ at its $k$th road intersection $v_{b,k}$. Upon arriving at its $(k\!-\!1)$th road intersection at time $t_{b,k-1}\!=\!t_{\text{sys}}$, the route ahead is denoted as
@@
To mitigate traffic congestion on each edge $(v_{b,k},v_{b,k+1})\!\in\!{P_{b}(t_{\text{sys}})}$ caused by CAVs, we define the control horizon for bus $b$ as the time interval $[t_{b,k-1},\hat{t}_{b,k}]$. The traffic flow on edge $(v_{b,k},v_{b,k+1})$ is monitored starting from time $t_{b,k-1}$, and CAVs are dynamically rerouted to avoid their entry onto the edge $(v_{b,k},v_{b,k+1})$ whenever congestion is detected. We make use of Fig. 3 for better illustration.
@@ V Simulation Study
This section presents simulation studies conducted on the Simulation of Urban MObility (SUMO) platform to evaluate the effectiveness and overall performance of the proposed routing approach in a mixed traffic network. We start by introducing the simulation setup and parameter configurations, and then present simulation results and analysis.
Figure 4: Visualization of the SUMO simulation environment, where the bus is depicted in yellow, CAVs are shown in red, and HVs are shown in blue. DLs are highlighted in grey, with bus stops marked in yellow.
TABLE I: Departure times of buses
Bus Index 1 2 3 4 5 6 7 8 9 10
Departure Time
[minute] 00 $6$ $12$ $18$ $24$ $30$ $36$ $42$ $48$ $54$
@@ V-A Simulation Setup
As shown in Fig. 4, we construct an urban road network in SUMO, which consists of $18$ road intersections and $5$ bus stops, with the road topology matching that shown in Fig. 1. We consider that buses follow a predefined route along the joint DLs, traveling from the origin stop (node $7$ in Fig. 1) to the destination stop (node $15$). All other edges in the network are designated as GPLs. We assume sensors are installed at each road intersection, allowing for real-time observation of the number of vehicles entering each edge within a predefined monitoring time window. Specifically, the monitoring time interval is set to $\Delta{T}\!=\!30$ seconds for joint DLs, and $\Delta\tilde{T}\!=\!60$ seconds for GPLs.
In the road network, we consider $10$ buses operating on the same route from the origin stop to the destination stop, with departures scheduled every $6$ minutes, as given in Table I. The scheduled on-time arrivals at bus stops are based on the free-flow travel time, with an added disturbance margin of $30$ seconds to account for potential delays caused by traffic signals at intersections between bus stops. CAVs and HVs enter the network at rates of $80$ and $120$ vehicles per minute, respectively. The OD pair of each CAV is set as $(7,15)$, while the OD pair for each HV is randomly selected from the set $\{(1,6),(7,15),(16,21)\}$.
@@
Figure 6: The dynamics of the accumulated CAV delay.
We compare the proposed coordinated routing approach with two baseline methods: Static Route Planning (SRP) without rerouting and Dynamic Route Planning (DRP) without rerouting. For ease of reference, we refer to the two methods as SRP and DRP without rerouting, respectively. In the SRP without rerouting approach, each CAV is assigned a shortest route based on the static free-flow travel times on the network edges and adheres to this route during its entire trip without any rerouting. In contrast, the DRP without rerouting incorporates real-time traffic data to update travel times on each edge and provides the shortest route based on the congestion conditions when the CAV enters the network. Similar to the SRP approach, no rerouting occurs for CAVs after deployment. The classic Dijkstra algorithm is employed to compute the shortest path for each CAV in the SRP without a rerouting scheme, while a modified Dijkstra algorithm using a weighted graph is applied to determine the shortest path in the DRP and proposed rerouting approaches. The weight graph is initially constructed based on free-flow travel time on each edge, and the weights are updated dynamically according to the predicted real-time travel conditions.
V-B Results and Analysis
We track the accumulated bus and CAV delays, measured as the total additional travel time incurred by all buses and CAVs relative to their free-flow travel times along each edge, over a 3500-second simulation horizon. The comparative results are presented in Fig. 5 and Fig. 6, respectively. As shown in Fig. 5, buses experience significantly greater delays under a static shortest-route assignment. In contrast, the proposed method significantly reduces bus delays by integrating real-time congestion estimation with efficient CAV rerouting. The results indicate that diverting CAVs away from congested edges, particularly those with bus stops during active dwell times, mitigates localized congestion along bus routes and improves schedule adherence. Similar trends are observed for CAV delays, as shown in Fig. 6. Although rerouting occasionally introduces detours for individual vehicles, the dynamic strategy yields lower overall delays. This improvement arises from proactively directing CAVs away from congestion-prone edges, especially those affected by stationary buses, thereby enhancing flow continuity throughout the network.
Figure 7: The accumulated bus delay at each stop.
@@ V-B Results and Analysis
Fig. 7 shows the accumulated trip delay for $10$ buses at each stop, comparing the DRP method without rerouting and the proposed approach. Bus stops indexed as $1$, $2$, $3$, and $4$ correspond to nodes $9$, $11$, $13$, and $15$ in Fig. 1. As is shown, the proposed method significantly reduces bus delays at all stops, achieving approximately a $53.6\%$ reduction in delay per stop compared to the DRP method. Fig. 8 further illustrates the accumulated delay experienced by each bus across all stops along the route. It is shown that our rerouting strategy consistently reduces delays for individual buses upon stop arrival, enhancing schedule adherence and overall transit reliability. Additionally, buses $5$ to $8$ experience relatively higher delays compared to buses $1$, $2$, $9$, and $10$, indicating that buses entering the network earlier or later encounter less congestion, aligning with the observed distribution of CAV and HV volumes in the transportation system.
Figure 9: Comparison of vehicle delays with and without using joint DLs.
To demonstrate the advantages of implementing joint DLs, Fig. 9 presents the average trip delays experienced by HVs and CAVs, comparing scenarios with and without the use of joint DLs. In the baseline case, all vehicles utilize GPLs exclusively. As illustrated in the figure, both HVs and CAVs benefit from substantially reduced trip delays under the joint DL configuration, saving around $38\%$ and $40\%$ trip times, respectively, indicating a marked enhancement in overall traffic performance and system-level efficiency.
The results presented above validate the effectiveness of the proposed rerouting strategy in preserving the transit priority of buses and enhancing their schedule adherence. By sharing joint DLs with buses, CAVs also benefit from the dynamic rerouting strategy through reduced trip delays. These outcomes demonstrate that the congestion-aware coordinated routing approach improves overall travel efficiency across the system.

### v2

@@ I Introduction
Connected and automated vehicles (CAVs) have emerged as a central element in the evolution of modern transportation systems. Advances in vehicle-to-vehicle, vehicle-to-infrastructure, and vehicle-to-everything communication technologies highlight the significant potential of CAVs in improving traffic efficiency, reducing energy consumption, and enhancing road safety [1, 2]. While the adoption of CAVs is projected to grow substantially over the coming decades, the transition towards a fully autonomous fleet is anticipated to unfold gradually [3, 4]. During this period, transportation networks will experience prolonged phases of mixed traffic, where CAVs and human-driven vehicles (HVs) coexist. Understanding how to enhance the traffic efficiency under increasing penetration rates of CAVs is therefore critical [5, 6, 7]. Extensive research and experimental studies have shown that heterogeneous traffic environments involving CAVs can significantly reduce travel times [8, 9, 10]. Among various approaches, implementing suitable right-of-way management is pivotal for ensuring the seamless integration of CAVs into existing transportation networks and for fully leveraging their benefits throughout this transition.
@@
To address the tradeoff inherent in implementing DLs for CAVs, a promising strategy has been proposed: the CAV and bus joint DL. Unlike conventional approaches that require constructing new lanes or displacing HVs, joint use of DLs by buses and CAVs repurposes existing DLs, which were originally designed for buses to ensure their on-time arrivals and maintain public transit efficiency [17]. While these lanes provide dedicated driving environments for transit operations, the reserved space cannot be taken by other non-transit vehicles, lowering the traffic efficiency [18]. Allowing CAVs to access these lanes can enhance lane utilization while maintaining transit reliability. In the existing literature, Luo et al. developed a framework for the optimal placement of CAV DLs to maximize benefits for all vehicles in the system [19]. Chen et al. explored a space-time modeling framework that demonstrates the mixed-use of DLs can enhance CAV travel efficiency and alleviate the overall traffic congestion [20]. A systematic framework investigating a variety of lane policies, such as exclusive automated vehicle lanes, regular vehicle lanes, and mixed-use lanes, was investigated in [12]. While the developed approaches provide new insights into the potential benefits of jointly using DLs, these methods remain limited in their ability to dynamically manage CAVs’ routing, particularly in coordinating vehicle behaviors across the network to balance traffic congestion reduction with maintaining reliable bus travel times, especially under fluctuating demand and a dynamic travel environment.
In this paper, we develop an efficient dynamic routing approach to coordinate CAVs traveling in a mixed traffic network, allowing for the DL sharing with buses. The goal is to prioritize the buses’ adherence to their scheduled timeliness while taking advantage of the joint DLs for CAVs to improve road capacity, reduce congestion, and enhance system-wide efficiency. By leveraging real-time traffic data, the proposed model enables estimation of potential tense flow on network edges and dynamically navigates the rerouting of CAVs to mitigate congestion on DLs. The main contributions of this work are as follows: (i) We model the rerouting problem of CAVs in a mixed traffic network where buses and CAVs jointly use DLs to fully take advantage of DLs while enhancing buses’ adherence to schedules. (ii) We incorporate a dynamic route condition estimation scheme in the rerouting framework, enabling real-time monitoring of traffic flow and timely adjustment of CAV routes when potential DL congestion is detected. Simulation performed in SUMO based on realistic traffic network data in San Francisco demonstrates the effectiveness of the proposed approach, indicating feasibility and significant improvements in traffic efficiency through coordinated use of joint DLs among buses and CAVs.
@@ II Problem Statement
Figure 1: An example illustrating the road transportation network, where bus stations ($\mathcal{S}$) are shown by red nodes, intersections ($\mathcal{I}$) are shown in black, and the dedicated lanes are shown by red edges.
Consider an urban road network represented by a directed graph $G\!=\!(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ denotes the set of nodes and $\mathcal{E}\!\subseteq\!\mathcal{V}\!\times\!\mathcal{V}$ denotes the set of directed edges. As illustrated in Fig. 1, the node set $\mathcal{V}$ is comprised of two subsets, denoted as $\mathcal{V}\!=\!\mathcal{S}\cup\mathcal{I}$, where $\mathcal{S}$ represents the set of bus stations and $\mathcal{I}$ represents the set of road intersections. Each directed edge $(v,v^{\prime})\in\mathcal{E}$ models a road segment from node $v$ to node $v^{\prime}$, with $v,v^{\prime}\in\mathcal{V}$. Consider each edge consisting of two lanes in the same direction, i.e., joint DL and GPL, where the GPL can be used by both CAVs and HVs, while DLs are reserved only for joint use by buses and CAVs.
Throughout this paper, we consider three types of vehicles traveling in the transportation network, including HVs, CAVs, and buses. The sets of HVs, CAVs, and buses within the monitoring time horizon are denoted by $\mathcal{N}^{\text{hv}}$, $\mathcal{N}^{\text{cav}}$, and $\mathcal{N}^{b}$, respectively. Accordingly, the set of all vehicles is represented as $\mathcal{N}\!=\!\mathcal{N}^{\text{hv}}\!\cup\!{\mathcal{N}^{\text{cav}}}\!\cup\!{\mathcal{N}^{b}}$. As illustrated in Fig. 2, buses use only DL and aim to arrive at each bus stop on time, following a specific timetable. In contrast, HVs can use only GPLs for travel purposes and route planning. With higher flexibility, CAVs can switch between joint DLs and GPLs to improve travel efficiency while minimizing interruptions to buses. Compared to employing only GPLs, driving on joint DLs enables CAVs to keep a smaller inter-vehicle distance, hence improving road capacity while avoiding disturbance caused by other HVs on GPLs.
@@ III-A Traffic Flow on Joint Dedicated Lane
[…]
@@ III-B Traffic Flow on General-Purpose Lane
[…]
@@
The traffic flow monitoring windows are defined by $\Delta{T}$ in (5) for joint DLs and by $\Delta{\tilde{T}}$ in (9) for GPLs, respectively, to capture real-time traffic conditions. A larger monitoring window puts greater emphasis on the macroscopic characteristics of traffic flow on network edges. With a different emphasis, the sizes of the two monitoring windows are not necessarily identical.
@@ IV-A Identification of CAVs for Rerouting
Let $\hat{t}_{b,k}$ denote the estimated arrival time of bus $b\!\in\!\mathcal{N}^{\text{b}}$ at its $k$th road intersection $v_{b,k}$. Upon arriving at its (k1)(k\!-\!1)th road intersection at time $t_{b,k-1}\!=\!t_{\text{sys}}$, the route ahead is denoted as
@@
To mitigate traffic congestion on each edge $(v_{b,k},v_{b,k+1})\!\in\!{P_{b}(t_{\text{sys}})}$ caused by CAVs, we define the control horizon for bus $b$ as the time interval $t_{b,k-1},\hat{t}_{b,k}]$. The traffic flow on edge $(v_{b,k},v_{b,k+1})$ is monitored starting from time $t_{b,k-1}$, and CAVs are dynamically rerouted to avoid their entry onto the edge $(v_{b,k},v_{b,k+1})$ whenever congestion is detected. We make use of Fig. 3 for better illustration.
@@ IV-B Route Optimization for Rerouted CAVs
Solving the above route optimization problem requires real-time traffic data collected from existing sensor infrastructures, such as loop detectors and cameras. These data are aggregated within monitoring windows to compute vehicles’ actual travel times on the current edges and to estimate their travel times on future edges. Incorporating travel time estimations enables long-horizon congestion prediction and route adjustment in advance, thus preventing vehicles from being trapped in severe congestion.
Remark 4
@@ V Simulation Study
This section presents simulation studies conducted on the Simulation of Urban MObility (SUMO) platform to evaluate the effectiveness and overall performance of the proposed approach in a mixed traffic network. We start by introducing the simulation setup and parameter configurations, and then present simulation results and analysis. The code implementation is available online. [^1] See code at: https://github.com/Tanlu-L/Coordinated-Routing-Approach-for-Enhancing-Bus-Timeliness-and-Travel-Efficiency
Figure 4: Illustration of the SUMO simulation environment, where the bus is depicted in yellow, CAVs are shown in red, and HVs are shown in blue. DLs are highlighted in grey, with bus stops marked in yellow.
TABLE I: Scheduled bus arrival times
Stop Station 1 Station 2 Station 3 Destination
Time [mm:ss] 01:03 03:46 07:39 09:27
@@ V-A Simulation Setup
As shown in Fig. 4, we construct an urban road network in SUMO, which consists of $18$ road intersections and $3$ bus stops, with the road topology matching that shown in Fig. 1. This network replicates Van Ness Avenue in San Francisco from Market Street to Lombard Street, which includes one lane of DL. The roadway geometry and signal settings were obtained from OpenStreetMap. We consider that buses follow a predefined route along the joint DLs, traveling from the origin (node $7$ in Fig. 1) to the destination (node $15$). All other edges in the network are designated as GPLs. We assume sensors are installed at road intersections, enabling real-time observation of the number of vehicles entering each edge within a predefined monitoring time window. The monitoring time interval is set to $\Delta{T}\!=\!30$ seconds for joint DLs, and $\Delta\tilde{T}\!=\!60$ seconds for GPLs.
In the road network, we modeled the workday peak-hour bus service with $10$ buses operating along the same route from the origin stop to the destination stop, departing at intervals of $6$ minutes. The scheduled on-time arrivals at bus stops were determined based on free-flow travel times, with an additional $60$ seconds of dwell time at each station and a $30$-second disturbance margin to account for potential delays caused by traffic signals between stops. The scheduled arrival times for a representative bus are provided in Table I. CAVs and HVs enter the network at rates of $8$ and $20$ vehicles per minute, respectively, corresponding to a $30\%$ CAV penetration rate. The OD pair for each CAV is set as $(7,15)$, while the OD pair for each HV is randomly selected from the set $\{(1,6),(7,15),(16,21)\}$.
We compare the proposed coordinated routing strategy against two baseline methods: (i) static route planning (SRP) without rerouting and (ii) dynamic route planning (DRP) with rerouting. In the SRP strategy, routes are assigned based on free-flow travel times and remain fixed throughout the trip. In the DRP approach, each CAV is assigned the shortest path derived from real-time traffic data at the entry. Then, rerouting is applied if a lower-travel-time path is detected within a 60-second monitoring window. In the proposed method, CAVs follow the route recommendation obtained by the predictive control approach described in Section IV. Different from our approach, congestion anticipation on future road segments is not incorporated in the SRP and DRP. For both SRP and DRP, a modified Dijkstra algorithm using a weighted graph is applied to determine the shortest path. The weighted graph is initially constructed based on free-flow travel time on each edge, and the weights are updated dynamically according to the real-time traffic conditions.
Regarding the car-following-model in all three scenarios, CAVs use the intelligent driver model with a short headway ($0.6$ s) to capture automation-enabled efficiency. The maximum speed is set to $72\,\text{km/h}$, consistent with corridor speed limits. HVs and buses use the Krauss model: HVs with moderate headways (about $1.3$ s), variability to reflect human drivers, and a maximum speed of $72\,\text{km/h}$. Buses operate with larger headways ($1.6$ s) and a lower maximum speed of $65\,\text{km/h}$. This configuration reflects the capacity benefits of CAVs while realistically modeling human drivers and bus operations.
V-B Results and Analysis
@@
TABLE II: Bus on-time performance and delays at each station.
Station 1 Station 2 Station 3
On-time percentage [%]
   SRP using joint DL 10 30 30
   DRP using joint DL 50 60 60
   Proposed method 90 90 90
We evaluate the accumulated bus delay, defined as the total additional travel time experienced by all buses relative to their scheduled arrival times upon arrival at the destination, over a 3,600-second simulation horizon. The comparative results are presented in Fig. 5. As shown, buses experience significant delays under a static shortest-route assignment. The DRP approach reduces delays to some extent, but still, buses experience noticeable delays. In contrast, the proposed method significantly reduces bus delays nearly back to zero by integrating real-time traffic conditions and potential congestion estimation with efficient CAV rerouting. The results indicate that diverting CAVs away from congested edges, particularly those with bus stops during active dwell times, mitigates localized congestion along bus routes and improves schedule adherence. Table II reports the on-time percentage of $10$ buses at each station. The on-time percentage is defined as the fraction of buses arriving within $60$ seconds of their scheduled arrival times as given in Table I. It shows that the on-time percentages of the SRP, DRP, and our proposed method are approximately $23\%$, $57\%$, and $90\%$, respectively, demonstrating the superior performance of our method.
Figure 6: The dynamics of the cumulative CAV travel time.
Figure 7: The dynamics of the cumulative HV travel time.
Figure 8: The average bus delay at each stop.
Fig. 6 and Fig. 7 show the total travel time of CAVs and HVs over the simulation horizon. The proposed method achieves the lowest overall travel time for both CAVs and HVs. This improvement arises from proactively directing CAVs from congestion-prone edges, especially those affected by dwelling buses, thereby enhancing flow continuity throughout the network. Note that the DRP approach results in longer travel times for CAVs compared to both the SRP and the proposed method. Although for HVs, DRP initially reduces travel times, after $1800$ seconds, the travel times increase and even exceed those under SRP. These observations are common in transportation routing problems, as CAVs simultaneously identify and switch to shorter routes, which in turn leads to new and often more severe congestion on those routes.
@@
Fig. 8 shows the average trip delay for $10$ buses at each station, compared between the SRP method using the joint DL and the proposed approach. Bus stations indexed as $1$, 22, and $3$ correspond to nodes $9$, $11$, and $13$ in Fig. 1. As we can observe, the proposed method reduces bus delays significantly at all stations, achieving an average reduction of about $90.7\%$ per station compared with the SRP method. Fig. 9 further depicts the accumulated delay experienced by each bus across all stations along the route, showing that our rerouting strategy consistently reduces delays for individual buses at stop arrivals, thereby improving schedule adherence and overall transit reliability. Note that bus $6$ experiences relatively higher delays than others, mainly due to the great number of vehicles in the system. This result indicates that buses entering the network earlier or later experience less congestion. Such a pattern also aligns with the observed distribution of CAVs and HVs volumes in the system.
Figure 10: The average trip delay under different CAV penetration rates.
We further evaluate the efficiency of the proposed method under different CAV penetration rates. Fig. 10 presents the average trip delays experienced by buses, HVs, and CAVs under penetration levels of $10\%$, $30\%$, and $50\%$, with the total travel demand set to $2,000$ vehicles per hour. The results show that HV delays steadily decrease as CAV penetration increases, while CAV and bus delays exhibit little variation between $30\%$ and $50\%$ penetration. These findings show that the benefits of increased penetration saturate beyond a certain threshold, with limited additional travel time savings.
Figure 11: 90th-percentile travel time comparison with and without joint DL.
To further evaluate the efficiency of the proposed method, we introduce a scenario termed SRP without using joint DL. In this scenario, we replicate the setting of a real dedicated bus lane by allowing only buses to use the lane, thereby ensuring that buses experience no delay, while CAVs follow their initially assigned shortest routes. The results show that allowing CAVs to share the joint DL reduces travel times for both CAVs and HVs, but at the expense of significantly increasing bus travel times. In contrast, the proposed method maintains bus travel times close to those in the dedicated bus lane scenario, while further reducing CAV and HV travel times compared to SRP with joint DL.
Given the above results, we validate the effectiveness of the proposed rerouting strategy in preserving bus transit priority and improving schedule adherence. Through the joint use of DLs with buses, both CAVs and HVs benefit from the dynamic rerouting with reduced trip delays. These findings demonstrate that the congestion-aware coordinated routing approach improves travel efficiency across the entire system.